Please help? Thanks.
2y(4-x)= x/2 where x=2

Answer by huntingmaster325(15)

2Y(4-x)=x/2 and x=2 plug in the 2 in place of the x
2Y(4-2)=2/2 first do what in the fraction.
2y*2=2\2 now do dividing and multipying
4y=1 now simpify it by dividing 4 from each side
answer is y=1/4